OpenStudy (anonymous):

find the slope: A. (-2,6), (-4, -7) B. y = (1/4)x + 2 C. (-4,-4), (-7, 24) HELP!!!!!!!!!!

OpenStudy (anonymous):

To find the slope using just cordinates you use this: \[\frac{ y ^{1}-y ^{2} }{ x ^{1}-x ^{2} }\]

OpenStudy (anonymous):

So, to do A and C you would start out with your coordinates (I will use A) and input them in the equation. If you do that then your equation should be this: \[\frac{ 6-(-7) }{ -2-(-4) }\]

OpenStudy (anonymous):

Subtracting a negative = adding a positive, so it would be: \[\frac{ 6+7 }{ -2+4 }\]

OpenStudy (anonymous):

Once you solve that, then that is your slope. I hope you'd know how to solve that. Just add the 6 and 7, and the -2 and 4. Then divide it. so 13 divided by 2, which is 6.5. That's your slope.

OpenStudy (anonymous):

For one like B, it's already in equation form, slope intercept form. Your slope is: y=mx+b. x= x intercept b= y intercept m= slope. Using that information what do you think the slope is?